# SonicID: User Identification on Smart Glasses with Acoustic Sensing

### v1

@@ Abstract.
Smart glasses have become more prevalent as they provide an increasing number of applications for users. They store various types of private information or can access it via connections established with other devices. Therefore, there is a growing need for user identification on smart glasses. In this paper, we introduce a low-power and minimally-obtrusive system called SonicID, designed to authenticate users on glasses. SonicID extracts unique biometric information from users by scanning their faces with ultrasonic waves and utilizes this information to distinguish between different users, powered by a customized binary classifier with the ResNet-18 architecture. SonicID can authenticate users within 0.12 seconds, with an energy consumption of 19.8 mAs per trial. A user study involving 24 participants confirms that SonicID achieves a true positive rate of 96.5%, a false positive rate of 4.1%, and a balanced accuracy of 96.2% using just 4 minutes of training data collected for each new user. This performance is relatively consistent across different remounting sessions and days. Given this promising performance, we further discuss the potential applications of SonicID and methods to improve its performance in the future.
@@
copyright: nonecopyright: acmcopyrightjournal: IMWUTjournalyear: 2024journalvolume: 0journalnumber: 0publicationmonth: 0price: 15.00doi: xxxxxx/xxxxxxx
@@ 1. Introduction
However, it is difficult to directly apply traditional frontal-camera-based authentication techniques, e.g. FaceID, on glasses because they require placing a camera in front of the user to capture their face. Researchers have put efforts into deploying cameras on glasses to capture biometric information of the user from their facial contours (Lim et al., 2023), eye movements (Zhang et al., 2018), or iris features (Li and Po-Jen, 2017). Though these camera-based methods achieve satisfactory performance of user authentication on glasses, they require instrumenting a camera which is relatively large compared to the small size of glasses and might partially block the user’s view. Furthermore, the cameras on commodity smart glasses, though can be smaller in size, usually point forwards to capture the environmental information. Barriers exist if they are to be directly used to capture the biometric information of the user’s face. Password-based authentication methods have been explored on smart glasses as well, especially with voice input (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). In addition, behavioral biometrics, such as tapping gestures (Islam et al., 2018), touch gestures (Jagmohan et al., 2016; Peng et al., 2017), and voice commands (Peng et al., 2017), have been leveraged to authenticate users. These methods require the user’s interaction and are notably vulnerable to shoulder-surfing attacks since these gestures are entered on a touchpad and the vocalized commands can be overheard by bystanders.
@@
Specifically, the speakers on the glasses emit encoded inaudible Frequency Modulated Continuous Waves (FMCW) towards the user’s face after they put on the glasses. The signals continuously and repeatedly sweep at different frequency ranges on two sides to avoid interference. The transmitted signals are reflected by the user’s face and captured by the microphones on the glasses. After signal processing using the received signals and transmitted signals, SonicID obtains unique acoustic patterns related to this user’s face. Considering everyone has different shapes of faces, the user’s biometric information is included in the captured acoustic patterns which can be learned to distinguish this user from other users and thus authenticate them securely. In SonicID, we feed the processed acoustic patterns into a customized binary classifier using the ResNet-18 architecture to extract distinct features so as to authenticate users. The classifier is trained with the training data collected when a user registers on the smart glasses and it keeps authenticating users in later usage of the device just like traditional user authentication techniques.
To validate the performance of our SonicID system, we conducted a user study with 24 participants. The study results demonstrate that a new user needs to provide 4 minutes of data to training the binary classifier on the first day when they use the smart glasses, to achieve a true positive rate of 96.5%, a false positive rate of 4.1% and a balanced accuracy of 96.2%. In this study, we also asked all participants to come back to test the performance of the system on two other different days. The results show that the performance of the system remains relatively consistent across three different days, especially for the false positive rate. If we continue to reduce the amount of training data to 1 minute for each participant, a true positive rate of 90.3%, a false positive rate of 1.6% and a balanced accuracy of 94.3% are achieved on Day 1. This study validates that SonicID achieves a satisfactory and stable authentication performance on smart glasses. Compared with frontal-camera-based authentication techniques such as FaceID, SonicID does not need a camera to be placed in front of the user. Other traditional authentication methods which take fingerprints or passwords as input can eliminate the need for a frontal camera. However, they require users’ operations for authentication while SonicID automatically logs valid users in when they put the device on. Moreover, SonicID is capable of successfully authenticating users within 0.12 seconds with a energy consumption of 19.8 mAs every time it authenticates the user. This low-power feature guarantees that SonicID can work approximately 28,000 times theoretically with the battery of Meta Ray-Ban (154 mAh) if no other operations are running on the device.
@@
  • We prototyped the system with one pair of speaker and microphone on each side of the glasses. The system maintains low-power, consuming only 19.8 mAs of energy every time it authenticates the user.
  • We conducted a user study with 24 participants, validating that only 4 minutes of training data for each new user guarantees that the system achieves a true positive rate of 96.5%, a false positive rate of 4.1% and a balanced accuracy of 96.2%. The performance remains relatively consistent in evaluations across different days.
@@ 2.1. Authentication on Non-Wearable Devices
Biometric authentication has gained popularity due to its perceived security and ease of use. Fingerprint sensing is widely used in current devices (Support, [n.d.]b). Face recognition technology, although increasingly common, can be compromised using photographs or videos of the user (Duc and Minh, 2009). To counter this, systems like Apple’s Face ID employ 3D reconstruction of facial features (Support, [n.d.]a). Iris scans offer another biometric alternative, noted for their uniqueness and difficulty to replicate (Daugman, 1994). Additionally, multimodal biometric systems have been explored, combining different physical attributes for enhanced security. For instance, Rokita et al. (Rokita et al., 2008) integrates face and hand images, while Marcel et al. (Marcel et al., 2010) combines face and speech recognition. EchoPrint (Zhou et al., 2018) merges facial recognition with acoustic analysis to counter spoofing attempts. VoiceLive (Zhang et al., 2016) and VoiceGesture (Zhang et al., 2017) utilize unique vocal characteristics and articulatory gestures, respectively, to thwart replay attacks. AirSign (Shao et al., 2020) innovatively employs acoustic sensing and motion sensors for air signature-based authentication.
@@
SonicID Glasses Acoustics Face No Yes Yes 24 96.2%
@@
2.2. Authentication on Wearable Devices
2.2.1. Smart Glasses
To enhance the security of smart glasses, a variety of novel authentication methods have been investigated. Traditional PIN-based mechanisms are notably vulnerable to shoulder-surfing attacks (Tari et al., 2006). This issue is exacerbated in smart glasses, where the PIN is entered on a touchpad, making it more visible to bystanders (Islam et al., 2018; Chan et al., 2015). Alternative approaches such as voice-based PIN entry have been explored (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). These methods employ a cipher to map the PIN to random digits, obscuring the password from eavesdroppers. However, this technique necessitates mental computations from users, potentially diminishing usability (Bailey et al., 2014). Camera-based authentication methods have also been a focus of research. The Glass OTP (Chan et al., 2015) system utilizes the camera in Google Glasses to scan a QR code on the user’s smartphone for authentication, although this method requires the user to carry a smartphone. C-Auth (Lim et al., 2023), which uses a downward-facing camera in the glasses to capture facial contours for authentication. Zhang et al. (Zhang et al., 2018) developed a continuous authentication system based on eye movement response to visual stimuli, detected by a camera. Another approach involves iris recognition, where internal infrared cameras are used for authentication (Li and Po-Jen, 2017). Behavioral biometrics have also been explored for smart glass authentication. Jagmohan et al. (Jagmohan et al., 2016) demonstrated a gesture-based continuous authentication system. The GlassGuard system (Peng et al., 2017), utilizes behavioral biometrics derived from touch gestures and voice commands for continuous authentication. Kawasaki et al. (Kawasaki and Sugiura, 2022) introduced an authentication method by observing the skin deformation during blinking, employing a photoreflector to measure blinks. Isobe et al. (Isobe and Murao, 2023) introduced an innovative approach using active acoustic sensing. This method involves transmitting acoustic signals through the nose using speakers integrated into the nose pads of the glasses, with the received signals captured by microphones. The acoustic structure of the nose serves as a biometric identifier, though this technique can be affected by the dryness of the nose. Furthermore, this prototype was not tested across different days so the stability of the system is unclear.
@@ 2.2.2. Other Wearable Devices
In the domain of smartwatches, researchers have extensively explored biometric-based authentication methods. For instance, Cornelius et al. (Cornelius et al., 2014) investigated the use of on-wrist bioimpedance responses for user authentication. Zhao et al. (Zhao et al., 2020) employed photoplethysmography signals for continuous authentication. Another innovative approach by Watanabe et al. (Watanabe et al., 2021) utilized active acoustic sensing, requiring users to perform four distinct hand poses for authentication. Further, Lee et al. (Lee et al., 2021) proposed a method leveraging vibration responses, measured through accelerometer and gyroscope sensors, to authenticate users. Recent advancements have introduced more sophisticated techniques. WristAcoustic (Huh et al., 2023), constructs a classifier based on three hand poses, utilizing a cue signal emitted from a surface transducer and recorded by a contact microphone. Additionally, WristConduct, (Sehrt et al., 2022), innovatively uses sound waves transmitted through the wrist bones, employing a bone conduction speaker and a laryngophone for authentication purposes.
@@
In addition to on-device authentication methods, researchers have explored approaches that incorporate secondary devices to facilitate user authentication. Nymi band (Nym, [n.d.]) is a wrist-worn wearable device that integrates fingerprint recognition, serving as an authentication mechanism for various devices and applications. Furthermore, ear-based authentication systems such as EarAE (Wu and He, 2023) and EarEcho (Gao et al., 2019) utilize acoustic sensing techniques to capture the unique structure of the ear canal, thereby offering a biometric signature for authentication purposes. ToothSonic (Wang et al., 2022), employs an earable-based system that utilizes toothprint-induced sonic waves for user verification. Similarly, Amesaka et al. (Amesaka et al., 2023) have demonstrated the potential of using sound leakage signals from hearables to capture the acoustic characteristics of the ear canal, auricle, or hand, which can then be employed for authentication. EarDynamic (Wang et al., 2021) leverages acoustic sensing in earables to assess both the static and dynamic motions of the ear canal during speech, providing a novel method for authentication.
2.3. Comparison between SonicID and Other Authentication Methods on Glasses
@@
To better compare our work with prior work of user authentication on glasses, we list all the projects that are most related to SonicID in Tab. 1. As shown in the table, SonicID does not need users’ interaction to perform authentication and achieves comparable balanced accuracy, if not better, to prior work on glasses. This performance is evaluated with a valid number of participants and remains relatively consistent across different remounting sessions and days. In the meantime, SonicID maintains low-power and minimally-obtrusive due to the advantages of acoustic sensors. In summary, SonicID contributes to the field by presenting the first acoustic-based method on smart glasses that scans users’ faces to obtain biometric information for user identification.
@@ 3.1. Face-based Biometric Authentication
Biometrics are unqiue physical characteristics that can be used for personal recognition (of Homeland Security, 2024). Face, as one of the most important parts of human body, has been a crucial source from which biometric information is extracted because the geometries of the facial surfaces of different people are different. Extensive research has been carried out to utilize this biometric information from the user’s face for authentication, especially captured by frontal cameras (Beumier and Acheroy, 2000; Bicego et al., 2006). One of the most commonly used authentication methods based on facial biometrics captured by cameras is Apple’s FaceID (Apple, 2024). It delivers impressive performance in authentication and has advantages of being more convenient and secure compared with password-based methods. The core sensing pricinple of SonicID is similar to FaceID. It instruments acoustic sensors on smart glasses to scan the user’s face in 3D with ultrasonic sound waves instead of infrared cameras and uses the obtained biometric information for authentication. The subsequent subsections introduce the principle of operation of SonicID in detail.

### v3

@@ Abstract.
Smart glasses have become more prevalent as they provide an increasing number of applications for users. They store various types of private information or can access it via connections established with other devices. Therefore, there is a growing need for user identification on smart glasses. In this paper, we introduce a low-power and minimally-obtrusive system called SonicID, designed to authenticate users on glasses. SonicID extracts unique biometric information from users by scanning their faces with ultrasonic waves and utilizes this information to distinguish between different users, powered by a customized binary classifier with the ResNet-18 architecture. SonicID can authenticate users by scanning their face for 0.06 seconds. A user study involving 40 participants confirms that SonicID achieves a true positive rate of 97.4%, a false positive rate of 4.3%, and a balanced accuracy of 96.6% using just 1 minute of training data collected for each new user. This performance is relatively consistent across different remounting sessions and days. Given this promising performance, we further discuss the potential applications of SonicID and methods to improve its performance in the future.
@@
ccs: Human-centered computing Ubiquitous and mobile devicesccs: Hardware Power and energycopyright: acmlicensedjournal: IMWUTjournalyear: 2024journalvolume: 8journalnumber: 4article: 169publicationmonth: 12doi: 10.1145/3699734
@@ 1. Introduction
However, it is difficult to directly apply traditional frontal-camera-based authentication techniques, e.g. FaceID, on glasses because they require placing a camera in front of the user to capture their face. Researchers have put efforts into deploying cameras on glasses to capture biometric information of the user from their facial contours (Lim et al., 2023), eye movements (Zhang et al., 2018), or iris features (Li and Po-Jen, 2017). Though these camera-based methods achieve satisfactory performance of user authentication on glasses, they require the placement of a camera which is relatively large compared to the small size of glasses and might partially block the user’s view. Furthermore, the cameras on commodity smart glasses, though can be smaller in size, usually point forwards to capture the environmental information. Barriers exist if they are to be directly used to capture the biometric information of the user’s face. Password-based authentication methods have been explored on smart glasses as well, especially with voice input (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). In addition, behavioral biometrics, such as tapping gestures (Islam et al., 2018), touch gestures (Jagmohan et al., 2016; Peng et al., 2017), and voice commands (Peng et al., 2017), have been leveraged to authenticate users. These methods require the user’s interaction and are notably vulnerable to shoulder-surfing attacks since these gestures are entered on a touchpad and the vocalized commands can be overheard by bystanders.
@@
Specifically, the speakers on the glasses emit encoded inaudible Frequency Modulated Continuous Waves (FMCW) towards the user’s face after they put on the glasses. The signals continuously and repeatedly sweep at different frequency ranges on two sides to avoid interference. The transmitted signals are reflected by the user’s face and captured by the microphones on the glasses. After signal processing using the received signals and transmitted signals, SonicID obtains unique acoustic patterns related to this user’s face. Considering that everyone has different facial shapes, the user’s biometric information is included in the captured acoustic patterns which can be learned to distinguish this user from other users and thus authenticate them securely. In SonicID, we feed the processed acoustic patterns into a customized binary classifier using the ResNet-18 architecture to extract distinct features so as to authenticate users. The classifier is trained with the training data collected when a user registers on the smart glasses and it keeps authenticating users in later usage of the device just like traditional user authentication techniques.
To validate the performance of our SonicID system, we conducted a user study with 40 participants. The study results demonstrate that a new user needs to provide 1 minute of data to train the binary classifier on the first day of using the smart glasses, to achieve a true positive rate of 97.4%, a false positive rate of 4.3%, and a balanced accuracy of 96.6%. In this study, we also asked all participants to come back to test the performance of the system on two other different days. The results show that the performance of the system remains relatively consistent across three different days, especially for the false positive rate. If 15 seconds of data is collected on a second day to fine-tune the trained model, the authentication performance of SonicID further improves across days, reaching a balanced accuracy around 90% on Day 2 and Day 3. This study validates that SonicID achieves a satisfactory and stable authentication performance on smart glasses. Compared with frontal-camera-based authentication techniques such as FaceID, SonicID does not need a camera to be placed in front of the user. Other traditional authentication methods which take fingerprints or passwords as input can eliminate the need for a frontal camera. However, they require users’ operations for authentication while SonicID automatically logs valid users in when they put the device on. Moreover, SonicID is capable of successfully authenticating users by scanning their face for only 0.06 seconds. Each time it authenticates users, SonicID, excluding the machine learning model inference part, consumes energy as little as 32.7 mJ. This low-power feature guarantees that SonicID can work approximately 31,500 times theoretically with the battery of Meta Ray-Ban (154 mAh) if the model inference of SonicID is implemented in real-time on a low-power chip such as MAX78002 (Devices, 2024) in the future work and no other operations are running on the device.
@@
  • We prototyped the system with one pair of speaker and microphone on each side of the glasses. The system maintains low-power, consuming only 32.7 mJ of energy every time it authenticates the user, excluding the machine learning model inference part.
  • We conducted a user study with 40 participants, validating that only 1 minute of training data for each new user guarantees that the system achieves a true positive rate of 97.4%, a false positive rate of 4.3%, and a balanced accuracy of 96.6%. The performance remains relatively consistent in evaluations across different days.
@@ 2.1. Authentication on Non-wearable Devices
Biometric authentication has gained popularity due to its perceived security and ease of use. Fingerprint sensing is widely used in current devices (Support, 2024b). Face recognition technologies, although increasingly common, can be compromised using photographs or videos of the user (Duc and Minh, 2009). To counter this, systems like Apple’s FaceID employ 3D reconstruction of facial features (Support, 2024a). Iris scans offer another biometric alternative, noted for their uniqueness and difficulty to replicate (Daugman, 1994). Additionally, multimodal biometric systems have been explored, combining different physical attributes for enhanced security. For instance, Rokita et al. (Rokita et al., 2008) integrates face and hand images, while Marcel et al. (Marcel et al., 2010) combines face and speech recognition. EchoPrint (Zhou et al., 2018) merges facial recognition with acoustic analysis to counter spoofing attempts. VoiceLive (Zhang et al., 2016) and VoiceGesture (Zhang et al., 2017) utilize unique vocal characteristics and articulatory gestures, respectively, to thwart replay attacks. AirSign (Shao et al., 2020) innovatively employs acoustic sensing and motion sensors for air signature-based authentication.
@@
2.2. Authentication on Wearable Devices
2.2.1. Smart Glasses
To enhance the security of smart glasses, a variety of novel authentication methods have been investigated. Traditional PIN-based mechanisms are notably vulnerable to shoulder-surfing attacks (Tari et al., 2006). This issue is exacerbated in smart glasses, where the PIN is entered on a touchpad, making it more visible to bystanders (Islam et al., 2018; Chan et al., 2015). Alternative approaches such as voice-based PIN entry have been explored (Bailey et al., 2014; Yadav et al., 2015; Li et al., 2021). These methods employ a cipher to map the PIN to random digits, obscuring the password from eavesdroppers. However, this technique necessitates mental computations from users, potentially diminishing usability (Bailey et al., 2014). Camera-based authentication methods have also been a focus of research. The Glass OTP (Chan et al., 2015) system utilized the camera in Google Glasses to scan a QR code on the user’s smartphone for authentication, although this method required the user to carry a smartphone. C-Auth (Lim et al., 2023) used a downward-facing camera in the glasses to capture facial contours for authentication. Zhang et al. (Zhang et al., 2018) developed a continuous authentication system based on eye movement response to visual stimuli, detected by a camera. Another approach involved iris recognition, where internal infrared cameras were used for authentication (Li and Po-Jen, 2017). Behavioral biometrics have also been explored for smart glass authentication. Jagmohan et al. (Jagmohan et al., 2016) demonstrated a gesture-based continuous authentication system. The GlassGuard system (Peng et al., 2017), utilized behavioral biometrics derived from touch gestures and voice commands for continuous authentication. Kawasaki et al. (Kawasaki and Sugiura, 2022) introduced an authentication method by observing the skin deformation during blinking, employing a photoreflector to measure blinks. Isobe et al. (Isobe and Murao, 2023) introduced an innovative approach using active acoustic sensing. This method involved transmitting acoustic signals through the nose using speakers integrated into the nose pads of the glasses, with the received signals captured by microphones. The acoustic structure of the nose served as a biometric identifier, though this technique could be affected by the dryness of the nose. Furthermore, this prototype was not tested across different days so the stability of the system was unclear.
@@ 2.2.1. Smart Glasses
SonicID Glasses Acoustics Face No Yes Yes 40 96.6%
@@ 2.2.2. Other Wearable Devices
In the domain of smartwatches, researchers have extensively explored biometric-based authentication methods. For instance, Cornelius et al. (Cornelius et al., 2014) investigated the use of on-wrist bioimpedance responses for user authentication. Zhao et al. (Zhao et al., 2020) employed photoplethysmography signals for continuous authentication. Another innovative approach by Watanabe et al. (Watanabe et al., 2021) utilized active acoustic sensing, requiring users to perform four distinct hand poses for authentication. Further, Lee et al. (Lee et al., 2021) proposed a method leveraging vibration responses, measured through accelerometer and gyroscope sensors, to authenticate users. Recent advancements have introduced more sophisticated techniques. WristAcoustic (Huh et al., 2023) constructed a classifier based on three hand poses, utilizing a cue signal emitted from a surface transducer and recorded by a contact microphone. Additionally, WristConduct (Sehrt et al., 2022) innovatively used sound waves transmitted through the wrist bones, employing a bone conduction speaker and a laryngophone for authentication purposes.
@@
In addition to on-device authentication methods, researchers have explored approaches that incorporate secondary devices to facilitate user authentication. Nymi band (Nym, [n. d.]) is a wrist-worn wearable device that integrates fingerprint recognition, serving as an authentication mechanism for various devices and applications. Furthermore, ear-based authentication systems such as EarAE (Wu and He, 2023) and EarEcho (Gao et al., 2019) utilized acoustic sensing techniques to capture the unique structure of the ear canal, thereby offering a biometric signature for authentication purposes. ToothSonic (Wang et al., 2022), employed an earable-based system that utilized toothprint-induced sonic waves for user verification. Similarly, Amesaka et al. (Amesaka et al., 2023) have demonstrated the potential of using sound leakage signals from hearables to capture the acoustic characteristics of the ear canal, auricle, or hand, which can then be employed for authentication. EarDynamic (Wang et al., 2021) leveraged acoustic sensing in earables to assess both the static and dynamic motions of the ear canal during speech, providing a novel method for authentication. F2Key (Duan et al., 2024) extracted acoustic features associated with the mouth movements when users spoke on a headphone and used the distinctive mapping between the acoustic features and the utterance for authentication.
2.3. Comparison between SonicID and Other Authentication Methods on Wearables
@@
To better compare our work with prior work of user authentication on wearables, we list all the projects that are most related to SonicID in Tab. 1. As shown in the table, SonicID does not need users’ interaction to perform authentication and achieves comparable balanced accuracy, if not better, to prior work on glasses. This performance is evaluated with a valid number of participants and remains relatively consistent across different remounting sessions and days. In the meantime, SonicID maintains low-power and minimally-obtrusive due to the advantages of acoustic sensors. Specifically compared with existing vision-based authentication methods on glasses (Lim et al., 2023; Zhang et al., 2018; Li and Po-Jen, 2017), SonicID achieves better performance across different remounting sessions. Moreover, we believe that SonicID surpasses vision-based methods in terms of the sensor size, the power consumption and the resilience to environmental lighting conditions because acoustic sensors are known to be smaller, consume less power and be less sensitive to lighting conditions compared with cameras.
In addition to cameras, other sensors, e.g., photoreflectors and acoustic sensors, have been utilized to perform user authentication on glasses as well as earables, headphones and wristbands in prior work. They all achieve promising balanced accuracies but many of them require user interaction, such as blinking (Kawasaki and Sugiura, 2022), toothprint (Wang et al., 2022), mouth movements (Duan et al., 2024) and hand poses (Huh et al., 2023), for authentication, while SonicID automatically identifies users when the device is put on. EarEcho (Gao et al., 2019) also automatically scans users’ ear canal with acoustic signals for authentication. However, the system is deployed on a pair of earphones and it captures the structure of the ear canal as biometrics, which are different from SonicID that captures the shape of the face as biometrics on a pair of glasses.
Isobe et al. (Isobe and Murao, 2023) innovatively integrated acoustic sensors into the nose pads of a pair of glasses for user authentication. Though it reaches a promising performance within the same day, the system is easily affected by the nasal conditions because the acoustic sensors are in direct contact with users’ nose. The experiments in the paper indicated that the balanced accuracy of the system drops from 91.0% to 88.4% and 83.0% respectively if the participant’s nose is wet or stuffed with cotton. Instead, SonicID places acoustic sensors on the hinges of the glasses, where many commercial smart glasses place their speakers and microphones. This helps the system suffer less from the impact of the internal state of users’ body, which guarantees better stability in practice. In addition, the proposed system in (Isobe and Murao, 2023) operates at a sampling rate up to 96 kHz, which is not supported by every commodity device. By comparison, SonicID requires a sampling rate of 50 kHz. The narrower operating frequency band limits the amount of useful information our system is able to obtain but SonicID still achieves a better authentication performance while proposing a lower technical requirement. Compared with the proposed system in (Isobe and Murao, 2023), it is easier for SonicID to be integrated into existing commodity smart glasses in the future because of its sensor position and technical requirements. In summary, SonicID contributes to the field by presenting the first acoustic-based method on smart glasses that scans users’ faces to obtain biometric information for user identification.
@@ 3.1. Face-based Biometric Authentication
Biometrics are unqiue physical characteristics that can be used for personal recognition (of Homeland Security, 2024). Face, as one of the most important parts of human body, has been a crucial source from which biometric information is extracted because the geometries of the facial surfaces of different people are different. Extensive research has been carried out to utilize this biometric information from the user’s face for authentication, especially captured by frontal cameras (Beumier and Acheroy, 2000; Bicego et al., 2006). One of the most commonly used authentication methods based on facial biometrics captured by cameras is Apple’s FaceID (Apple, 2024). It delivers impressive performance in authentication and has advantages of being more convenient and secure compared with password-based methods. The core sensing pricinple of SonicID is similar to FaceID. It places acoustic sensors on smart glasses to scan the user’s face in 3D with ultrasonic sound waves instead of infrared cameras and uses the obtained biometric information for authentication.
To realize this goal, we decided to employ the calculation of echo profiles as the main data processing method because it exhibits promising performance in extracting patterns related to facial movements (Li et al., 2022, 2024a) and body movements (Mahmud et al., 2023) in prior work. Nevertheless, all these application scenarios are motion-related while our work aims to exploit echo profiles to represent the acoustic patterns of facial scanning and validate the feasibility of user identification based on these patterns. The main contribution of this paper is not to propose the echo profile calculation. Instead, we use echo profiles as a standard acoustic data processing technique, similar to mel-spectrograms and Doppler effects, and our goal is to explore how the acoustic patterns can be better exploited to serve the purpose of user identification. To reach this end, we designed several specific steps in the pipeline such as the selection of static instances (Sec. 3.3) and the two-stage training technique for the machine learning model (Sec. 3.4), to boost the authentication performance. Moreover, we experimented on the effectiveness of each step of the system pipeline including the choice of two pairs of speakers and microphones, the length and number of the static instances, and the pre-training stage in Sec. 6. The evaluation results validated the effectiveness of the data processing and machine learning pipeline of SonicID. The major contribution of SonicID is to propose the first acoustic-based method on smart glasses that is able to perform user identification by scanning users’ full face. The subsequent subsections introduce the principle of operation of SonicID in detail.